# A Tverberg-type problem of Kalai: Two negative answers to questions of Alon and Smorodinsky, and the power of disjointness

### v1

@@ Acknowledgement
As a participant in the ECOPRO Summer Student Research Program, Wenchong Chen acknowledges the organizers, especially Hong Liu, and the members of IBS for their support and hospitality. Zixiang Xu would like to thank Hong Liu and Chong Shangguan for their encouragement over the past few years to pursue topics in combinatorial convexity, in particular, he gratefully acknowledges Chong Shangguan’s gift of the print edition of [4] in 2023. While working on this project in September, Zixiang Xu thanks Jiangdong Ai, Fankang He, Yulai Ma, Shuaichao Wang, and Yongtang Shi for their hospitality at Nankai University; Jie Han for his hospitality at the Beijing Institute of Technology; and Luyi Li, Ke Ye, and Qiang Zhou for their hospitality at the Chinese Academy of Sciences.

### v2

@@ Acknowledgement
The authors would like to thank Noga Alon, Gil Kalai and Shakhar Smorodinsky for their helpful comments. In particular, they thank Shakhar Smorodinsky for bringing to their attention the result of [14], which significantly simplifies the proof of Theorem 1.6.
As a participant in the ECOPRO Summer Student Research Program, Wenchong Chen acknowledges the organizers, especially Hong Liu, and the members of IBS for their support and hospitality. Zixiang Xu would like to thank Hong Liu and Chong Shangguan for their encouragement over the past few years to pursue topics in combinatorial convexity, in particular, he gratefully acknowledges Chong Shangguan’s gift of the print edition of [4] in 2023. While working on this project in September, Zixiang Xu thanks Jiangdong Ai, Fankang He, Yulai Ma, Yongtang Shi and Shuaichao Wang for their hospitality at Nankai University; Jie Han for his hospitality at the Beijing Institute of Technology; and Luyi Li, Ke Ye, and Qiang Zhou for their hospitality at the Chinese Academy of Sciences.